OpenStudy (anonymous):

The first term of an arithmetic sequence is -5, and the tenth term is 13. Find the common difference. 8/9 1.8 2

hartnn (hartnn):

do you know the formula for n'th term of arithmetic sequence ?

OpenStudy (anonymous):

no wat u see is all i have

hartnn (hartnn):

its this : \(\large a_n=a_1+(n-1)d\) an is the n'th term, a1 = 1st term, here a1 = -5 d = common difference

hartnn (hartnn):

if we have 10th term, you just plug in n=10 and a10 is given = 13

hartnn (hartnn):

so put n=10 an =13 a1 = -5 in that formula! only "d" is unknown, can you find it ?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

no ? whats the problem you are facing ? you have the formula, and the values, did you try plugging in ?

OpenStudy (anonymous):

i hav tried but i just dont get it. math is my worst subject so yea

hartnn (hartnn):

we have 13 = -5 + (10-1)d yeah ? start by adding 5 on both sides, what do u get ?

OpenStudy (anonymous):

ummmm 17 and 0?

hartnn (hartnn):

13+5 =18

OpenStudy (anonymous):

oh whoops

hartnn (hartnn):

so, 18 = 9d just divide 9 on both sides, d= .. ?

OpenStudy (anonymous):

1?

hartnn (hartnn):

how?

OpenStudy (anonymous):

2?

hartnn (hartnn):

yes!

OpenStudy (anonymous):

thanx

hartnn (hartnn):

welcome ^_^